Gamma Industries has net income of $500,000, and it has 1,305,000 shares of common stock outstanding. The company's stock currently trades at $39 a share. Gamma is considering a plan in which it will use available cash to repurchase 20% of its shares in the open market at the current $39 stock price. The repurchase is expected to have no effect on net income or the company's P/E ratio. What will be its stock price following the stock repurchase? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Net income = $500,000

Shares outstanding = 1,305,000

Current market price (CMP) = $39

Hence,

EPS = Net income / Shares outstanding
EPS = 500,000 / 1,305,000
EPS = 0.3831

P/E = CMP / EPS
P/E = 39 / 0.3831
P/E = 101.80

Shares repurchased = 20% of shares outstanding
Shares repurchased = 20% of 1,305,000
Shares repurchased = 261,000

New shares outstanding = Shares outstanding - shares repurchased
New shares outstanding = 1,305,000 - 261,000
New shares outstanding = 1,044,000

New EPS = Net income / New outstanding share
New EPS = 500,000 / 1,044,000
New EPS = 0.4789

New stock price = P/E x New EPS
New stock price = 101.80 x 0.4789
New stock price = $48.75

Hence stock price following stock repurchase will be $48.75